\begin{thebibliography}{99}
%\cite{Bunce:1976yb}
\bibitem{Bunce:1976yb}
%  G.~Bunce, R.~Handler, R.~March, P.~Martin, L.~Pondrom, M.~Sheaff, K.~J.~Heller and O.~Overseth {\it et al.},
  G.~Bunce {\it et al.},
  %``Lambda0 Hyperon Polarization in Inclusive Production by 300-GeV Protons on Beryllium.,''
  Phys.\ Rev.\ Lett.\  {\bf 36}, 1113 (1976).
 %\cite{Adams:1991cs}
%\bibitem{Adams:1991cs}
  D.~L.~Adams {\it et al.} [E704 Collaboration],
  %``Analyzing power in inclusive pi+ and pi- production at high x(F) with a 200-GeV polarized proton beam,''
  Phys.\ Lett.\ B {\bf 264}, 462 (1991).
  %%CITATION = PHLTA,B264,462;%%
%\cite{:2008qb}
%\bibitem{:2008qb}
  B.~I.~Abelev {\it et al.} [STAR Collaboration],
  %``Forward Neutral Pion Transverse Single Spin Asymmetries in p+p Collisions at \sqrt{s}=200 GeV,''
  Phys.\ Rev.\ Lett.\  {\bf 101}, 222001 (2008).
%  [arXiv:0801.2990 [hep-ex]].
  %%CITATION = PRLTA,101,222001;%%
%\cite{Adler:2005in}
%\bibitem{Adler:2005in}
  S.~S.~Adler {\it et al.}  [PHENIX Collaboration],
  %``Measurement of transverse single-spin asymmetries for mid-rapidity production of neutral pions and charged hadrons in polarized p+p collisions at s**(1/2) = 200-GeV,''
  Phys.\ Rev.\ Lett.\  {\bf 95}, 202001 (2005).
%  [hep-ex/0507073].
  %%CITATION = HEP-EX/0507073;%%
%\cite{:2008mi}
%\bibitem{:2008mi}
  I.~Arsene {\it et al.}  [BRAHMS Collaboration],
  %``Single Transverse Spin Asymmetries of Identified Charged Hadrons in Polarized p+p Collisions at s**(1/2) = 62.4-GeV,''
  Phys.\ Rev.\ Lett.\  {\bf 101}, 042001 (2008).
%  [arXiv:0801.1078 [nucl-ex]].
  %%CITATION = ARXIV:0801.1078;%%

%\cite{Kane:1978nd}
\bibitem{Kane:1978nd}
  G.~L.~Kane, J.~Pumplin and W.~Repko,
  %``Transverse Quark Polarization in Large p(T) Reactions, e+ e- Jets, and Leptoproduction: A Test of QCD,''
  Phys.\ Rev.\ Lett.\  {\bf 41}, 1689 (1978).
  %%CITATION = PRLTA,41,1689;%%

 %\cite{Ma:2008gm}
\bibitem{Ma:2008gm}
  J.~P.~Ma and H.~Z.~Sang,
  %``Partonic State and Single Transverse Spin Asymmetry in Drell-Yan Process,''
  JHEP {\bf 0811}, 090 (2008).
 %[arXiv:0809.4811 [hep-ph]].
  %%CITATION = ARXIV:0809.4811;%%
  %\cite{Ma:2008cj}


  %\cite{Sivers:1990fh}
\bibitem{Sivers:1990fh}
  D.~W.~Sivers,
  %``Hard scattering scaling laws for single spin production asymmetries,''
  Phys.\ Rev.\ D {\bf 43}, 261 (1991).
  %%CITATION = PHRVA,D43,261;%%


  %\cite{Collins:1992kk}
\bibitem{Collins:1992kk}
  J.~C.~Collins,
  %``Fragmentation of transversely polarized quarks probed in transverse momentum distributions,''
  Nucl.\ Phys.\ B {\bf 396}, 161 (1993).
%  [hep-ph/9208213].
  %%CITATION = HEP-PH/9208213;%%


 %\cite{Efremov:1981sh}
\bibitem{Efremov:1981sh}
  A.~V.~Efremov and O.~V.~Teryaev,
  %``On Spin Effects In Quantum Chromodynamics,''
  Sov.\ J.\ Nucl.\ Phys.\  {\bf 36}, 140 (1982)
  [Yad.\ Fiz.\  {\bf 36}, 242 (1982)];
  %%CITATION = SJNCA,36,140;%%
%\cite{Efremov:1984ip}
%\bibitem{Efremov:1984ip}
  %``QCD Asymmetry And Polarized Hadron Structure Functions,''
  Phys.\ Lett.\ B {\bf 150}, 383 (1985).
  %%CITATION = PHLTA,B150,383;%%


%\cite{Qiu:1991pp}
\bibitem{Qiu:1991pp}
  J.-w.~Qiu and G.~F.~Sterman,
  %``Single transverse spin asymmetries,''
  Phys.\ Rev.\ Lett.\  {\bf 67}, 2264 (1991);
  %CITATION = PRLTA,67,2264;%%
%\cite{Qiu:1991wg}
%bibitem{Qiu:1991wg}
%  J.-w.~Qiu and G.~F.~Sterman,
  %``Single transverse spin asymmetries in direct photon production,''
  Nucl.\ Phys.\ B {\bf 378}, 52 (1992).
 %%CITATION = NUPHA,B378,52;%%

   %\cite{Ji:1992eu}
\bibitem{Ji:1992eu}
  X.~-D.~Ji,
  %``Gluon correlations in the transversely polarized nucleon,''
  Phys.\ Lett.\ B {\bf 289}, 137 (1992).
  %%CITATION = PHLTA,B289,137;%%
  %76 citations counted in INSPIRE as of 05 Mar 2013


    %\cite{Beppu:2010qn}
\bibitem{Beppu:2010qn}
  H.~Beppu, Y.~Koike, K.~Tanaka and S.~Yoshida,
  %``Contribution of Twist-3 Multi-Gluon Correlation Functions to Single Spin Asymmetry in Semi-Inclusive Deep Inelastic Scattering,''
  Phys.\ Rev.\ D {\bf 82}, 054005 (2010).
%  [arXiv:1007.2034 [hep-ph]].
  %%CITATION = ARXIV:1007.2034;%%
  %25 citations counted in INSPIRE as of 05 Mar 2013

  %\cite{Yuan:2009dw}
\bibitem{Yuan:2009dw}
  F.~Yuan and J.~Zhou,
  %``Collins Fragmentation and the Single Transverse Spin Asymmetry,''
  Phys.\ Rev.\ Lett.\  {\bf 103}, 052001 (2009).
 % [arXiv:0903.4680 [hep-ph]].
  %%CITATION = ARXIV:0903.4680;%%
  %29 citations counted in INSPIRE as of 01 Aug 2013


%\cite{Hoyer:2006hu}
\bibitem{Hoyer:2006hu}
  P.~Hoyer and M.~Jarvinen,
  %``Single spin asymmetry at large x(F) and k-perpendicular,''
  JHEP {\bf 0702}, 039 (2007).
%  [hep-ph/0611293];
  %%CITATION = HEP-PH/0611293;%%
%\cite{Hoyer:2008fp}
%\bibitem{Hoyer:2008fp}
  P.~Hoyer, M.~Jarvinen and S.~Kurki,
  %``Factorization at fixed Q**2(1-x),''
  JHEP {\bf 0810}, 086 (2008).
%  [arXiv:0808.0626 [hep-ph]].
  %%CITATION = ARXIV:0808.0626;%%

    %\cite{Kochelev:1996pv}
\bibitem{Kochelev:1996pv}
  N.~I.~Kochelev,
  %``Anomalous quark chromomagnetic moment induced by instantons,''
  Phys.\ Lett.\ B {\bf 426}, 149 (1998)
 .
  %%CITATION = HEP-PH/9610551;%%
  %41 citations counted in INSPIRE as of 08 Oct 2013
  %\cite{Kochelev:2013zoa}
%\bibitem{Kochelev:2013zoa}
  N.~Kochelev and N.~Korchagin,
  %``Anomalous Quark Chromomagnetic Moment and Single-Spin Asymmetries,''
  arXiv:1308.4857 [hep-ph].
  %%CITATION = ARXIV:1308.4857;%%

%\cite{Qian:2011ya}
\bibitem{Qian:2011ya}
  Y.~Qian and I.~Zahed,
  %``Single Spin Asymmetry through QCD Instantons,''
  Phys.\ Rev.\ D {\bf 86}, 014033 (2012)
  [Erratum-ibid.\ D {\bf 86}, 059902 (2012)].
  %%CITATION = ARXIV:1112.4552;%%
  %7 citations counted in INSPIRE as of 23 Aug 2013


  %\cite{Ahmedov:1999ne}
\bibitem{Ahmedov:1999ne}
  A.~Ahmedov, I.~V.~Akushevich, E.~A.~Kuraev and P.~G.~Ratcliffe,
  %``Single spin asymmetries for small angle pion production in high-energy hadron collisions,''
  Eur.\ Phys.\ J.\ C {\bf 11}, 703 (1999).
  %%CITATION = HEP-PH/9902418;%%
  %11 citations counted in INSPIRE as of 01 Aug 2013
  %\cite{Ahmedov:2002yh}
%\bibitem{Ahmedov:2002yh}
  A.~Ahmedov, E.~N.~Antonov, E.~Bartos, E.~A.~Kuraev and E.~Zemlyanaya,
  %``Single spin asymmetry in pion production in polarized proton proton collisions and odderon,''
  J.\ Phys.\ G {\bf 23}, 521 (2003)
  [J.\ Phys.\ G {\bf 29}, 521 (2003)].
  %%CITATION = HEP-PH/0207099;%%
  %1 citations counted in INSPIRE as of 01 Aug 2013

%\cite{Kovchegov:2012ga}
\bibitem{Kovchegov:2012ga}
  Y.~V.~Kovchegov and M.~D.~Sievert,
  %``A New Mechanism for Generating a Single Transverse Spin Asymmetry,''
  Phys.\ Rev.\ D {\bf 86}, 034028 (2012)
  [Erratum-ibid.\ D {\bf 86}, 079906 (2012)].
  %%CITATION = ARXIV:1201.5890;%%
  %11 citations counted in INSPIRE as of 23 Aug 2013

  %\cite{Bartels:1995iu}
\bibitem{Bartels:1995iu}
  J.~Bartels, B.~I.~Ermolaev and M.~G.~Ryskin,
  %``Nonsinglet contributions to the structure function g1 at small x,''
  Z.\ Phys.\ C {\bf 70}, 273 (1996);
  %%CITATION = HEP-PH/9507271;%%
  %131 citations counted in INSPIRE as of 01 Aug 2013
%\cite{Bartels:1996wc}
%\bibitem{Bartels:1996wc}
 % J.~Bartels, B.~I.~Ermolaev and M.~G.~Ryskin,
  %``Flavor singlet contribution to the structure function G(1) at small x,''
  Z.\ Phys.\ C {\bf 72}, 627 (1996).
  %%CITATION = HEP-PH/9603204;%%
  %125 citations counted in INSPIRE as of 01 Aug 2013

  %\cite{oai:arXiv.org:hep-ph/0202231}
\bibitem{oai:arXiv.org:hep-ph/0202231}
  P.~.Hagler, B.~Pire, L.~Szymanowski and O.~V.~Teryaev,
  %``Hunting the QCD odderon in hard diffractive electroproduction of two pions,''
  Phys.\ Lett.\ B {\bf 535}, 117 (2002)
  [Erratum-ibid.\ B {\bf 540}, 324 (2002)]
  [hep-ph/0202231].
  %%CITATION = HEP-PH/0202231;%%
  %66 citations counted in INSPIRE as of 07 Feb 2014

  %\cite{Boer:2006rj}
\bibitem{Boer:2006rj}
  D.~Boer, A.~Dumitru and A.~Hayashigaki,
  %``Single transverse-spin asymmetries in forward pion production at high energy: Incorporating small-x effects in the target,''
  Phys.\ Rev.\ D {\bf 74}, 074018 (2006).
  %%CITATION = HEP-PH/0609083;%%
  %16 citations counted in INSPIRE as of 01 Aug 2013
  %\cite{Kang:2011ni}
%\bibitem{Kang:2011ni}
  Z.~-B.~Kang and F.~Yuan,
  %``Single Spin Asymmetry Scaling in the Forward Rapidity Region at RHIC,''
  Phys.\ Rev.\ D {\bf 84}, 034019 (2011).
  %%CITATION = ARXIV:1106.1375;%%
  %11 citations counted in INSPIRE as of 01 Aug 2013


%\cite{Metz:2011wb}
\bibitem{Metz:2011wb}
  A.~Metz and J.~Zhou,
  %``Distribution of linearly polarized gluons inside a large nucleus,''
  Phys.\ Rev.\ D {\bf 84}, 051503 (2011).
  %%CITATION = ARXIV:1105.1991;%%
  %17 citations counted in INSPIRE as of 01 Aug 2013
%\cite{Schafer:2013mza}
%\bibitem{Schafer:2013mza}
  A.~Sch\"{a}fer and J.~Zhou,
  %``On the process dependent nuclear $k_\perp$ broadening effect,''
  arXiv:1305.5042 [hep-ph].
  %%CITATION = ARXIV:1305.5042;%%
  %\cite{Chen:2013zpy}
%\bibitem{Chen:2013zpy}
  L.~Chen, J.~-h.~Gao and Z.~-t.~Liang,
  %``Nuclear dependences of azimuthal asymmetries in the Drell-Yan process,''
  arXiv:1308.3746 [hep-ph].
  %%CITATION = ARXIV:1308.3746;%%



%\cite{Bartels:1980pe}
\bibitem{Bartels:1980pe}
  J.~Bartels,
  %``High-Energy Behavior in a Nonabelian Gauge Theory. 2. First Corrections to T(n--->m) Beyond the Leading LNS Approximation,''
  Nucl.\ Phys.\ B {\bf 175}, 365 (1980).
  %%CITATION = NUPHA,B175,365;%%
  %302 citations counted in INSPIRE as of 02 Aug 2013
%\cite{Kwiecinski:1980wb}
%\bibitem{Kwiecinski:1980wb}
  J.~Kwiecinski and M.~Praszalowicz,
  %``Three Gluon Integral Equation and Odd c Singlet Regge Singularities in QCD,''
  Phys.\ Lett.\ B {\bf 94}, 413 (1980).
  %%CITATION = PHLTA,B94,413;%%
  %249 citations counted in INSPIRE as of 02 Aug 2013

%\cite{Hatta:2005as}
\bibitem{Hatta:2005as}
  Y.~Hatta, E.~Iancu, K.~Itakura and L.~McLerran,
  %``Odderon in the color glass condensate,''
  Nucl.\ Phys.\ A {\bf 760}, 172 (2005).
  %%CITATION = HEP-PH/0501171;%%
  %43 citations counted in INSPIRE as of 02 Aug 2013


  %\cite{Kovchegov:2003dm}
\bibitem{Kovchegov:2003dm}
  Y.~V.~Kovchegov, L.~Szymanowski and S.~Wallon,
  %``Perturbative odderon in the dipole model,''
  Phys.\ Lett.\ B {\bf 586}, 267 (2004).
  %%CITATION = HEP-PH/0309281;%%
  %34 citations counted in INSPIRE as of 03 Aug 2013

  %\cite{Bartels:1999yt}
\bibitem{Bartels:1999yt}
  J.~Bartels, L.~N.~Lipatov and G.~P.~Vacca,
  %``A New odderon solution in perturbative QCD,''
  Phys.\ Lett.\ B {\bf 477}, 178 (2000).
  %%CITATION = HEP-PH/9912423;%%
  %120 citations counted in INSPIRE as of 05 Aug 2013

    %\cite{McLerran:1993ni}
\bibitem{McLerran:1993ni}
  L.~D.~McLerran and R.~Venugopalan,
  %``Computing quark and gluon distribution functions for very large nuclei,''
  Phys.\ Rev.\ D {\bf 49}, 2233 (1994);
  %%CITATION = HEP-PH/9309289;%%
  %1239 citations counted in INSPIRE as of 07 Aug 2013
  %\cite{McLerran:1993ka}
%\bibitem{McLerran:1993ka}
 % L.~D.~McLerran and R.~Venugopalan,
  %``Gluon distribution functions for very large nuclei at small transverse momentum,''
  Phys.\ Rev.\ D {\bf 49}, 3352 (1994).
  %%CITATION = HEP-PH/9311205;%%
  %913 citations counted in INSPIRE as of 07 Aug 2013


%\cite{Jeon:2004rk}
\bibitem{Jeon:2004rk}
  S.~Jeon and R.~Venugopalan,
  %``Random walks of partons in SU(N(c)) and classical representations of color charges in QCD at small x,''
  Phys.\ Rev.\ D {\bf 70}, 105012 (2004);
  %%CITATION = HEP-PH/0406169;%%
  %35 citations counted in INSPIRE as of 03 Aug 2013
  %\cite{Jeon:2005cf}
%\bibitem{Jeon:2005cf}
 % S.~Jeon and R.~Venugopalan,
  %``A Classical Odderon in QCD at high energies,''
  Phys.\ Rev.\ D {\bf 71}, 125003 (2005).
  %%CITATION = HEP-PH/0503219;%%
  %26 citations counted in INSPIRE as of 03 Aug 2013


  %\cite{Burkardt:2000za}
\bibitem{Burkardt:2000za}
  M.~Burkardt,
  %``Impact parameter dependent parton distributions and off forward parton distributions for zeta ---> 0,''
  Phys.\ Rev.\ D {\bf 62}, 071503 (2000)
  [Erratum-ibid.\ D {\bf 66}, 119903 (2002)];
  %%CITATION = HEP-PH/0005108;%%
  %383 citations counted in INSPIRE as of 13 Aug 2013
  %\cite{Burkardt:2002hr}
%\bibitem{Burkardt:2002hr}
 % M.~Burkardt,
  %``Impact parameter space interpretation for generalized parton distributions,''
  Int.\ J.\ Mod.\ Phys.\ A {\bf 18}, 173 (2003).
  %%CITATION = HEP-PH/0207047;%%
  %354 citations counted in INSPIRE as of 13 Aug 2013

  %\cite{Burkardt:2002ks}
\bibitem{Burkardt:2002ks}
  M.~Burkardt,
  %``Impact parameter dependent parton distributions and transverse single spin asymmetries,''
  Phys.\ Rev.\ D {\bf 66}, 114005 (2002).
  %%CITATION = HEP-PH/0209179;%%
  %123 citations counted in INSPIRE as of 15 Aug 2013


  %\cite{Gockeler:2006zu}
\bibitem{Gockeler:2006zu}
  M.~Gockeler {\it et al.}  [QCDSF and UKQCD Collaborations],
  %``Transverse spin structure of the nucleon from lattice QCD simulations,''
  Phys.\ Rev.\ Lett.\  {\bf 98}, 222001 (2007).
  %%CITATION = HEP-LAT/0612032;%%
  %102 citations counted in INSPIRE as of 09 Aug 2013

  %\cite{Albacete:2010sy}
\bibitem{Albacete:2010sy}
  J.~L.~Albacete, N.~Armesto, J.~G.~Milhano, P.~Quiroga-Arias and C.~A.~Salgado,
  %``AAMQS: A non-linear QCD analysis of new HERA data at small-x including heavy quarks,''
  Eur.\ Phys.\ J.\ C {\bf 71}, 1705 (2011).
    %%CITATION = ARXIV:1012.4408;%%
  %56 citations counted in INSPIRE as of 06 Aug 2013

  %\cite{Dumitru:2002qt}
\bibitem{Dumitru:2002qt}
  A.~Dumitru and J.~Jalilian-Marian,
  %``Forward quark jets from protons shattering the colored glass,''
  Phys.\ Rev.\ Lett.\  {\bf 89}, 022301 (2002).
   %%CITATION = HEP-PH/0204028;%%
  %111 citations counted in INSPIRE as of 06 Aug 2013

%\cite{Chirilli:2011km}
\bibitem{Chirilli:2011km}
  G.~A.~Chirilli, B.~-W.~Xiao and F.~Yuan,
  %``One-loop Factorization for Inclusive Hadron Production in $pA$ Collisions in the Saturation Formalism,''
  Phys.\ Rev.\ Lett.\  {\bf 108}, 122301 (2012);
  %%CITATION = ARXIV:1112.1061;%%
  %18 citations counted in INSPIRE as of 06 Aug 2013
  %\cite{Chirilli:2012jd}
%\bibitem{Chirilli:2012jd}
 % G.~A.~Chirilli, B.~-W.~Xiao and F.~Yuan,
  %``Inclusive Hadron Productions in pA Collisions,''
  Phys.\ Rev.\ D {\bf 86}, 054005 (2012).
  %%CITATION = ARXIV:1203.6139;%%
  %13 citations counted in INSPIRE as of 06 Aug 2013

  %\cite{Schafer:2013opa}
\bibitem{Schafer:2013opa}
  A.~Schafer and J.~Zhou,
  %``A note on the scale evolution of tri-gluon correlations,''
  arXiv:1308.4961 [hep-ph].
  %%CITATION = ARXIV:1308.4961;%%

%\cite{Nogach:2012sh}
\bibitem{Nogach:2012sh}
  L.~Nogach [AnDY Collaboration],
  %``Measurements of Forward Jet Production in Polarized $pp$ Collisions at $\sqrt{s}=500$ GeV,''
  arXiv:1212.3437 [hep-ex].
  %%CITATION = ARXIV:1212.3437;%%
  %7 citations counted in INSPIRE as of 06 Aug 2013
%\cite{Bland:2013pkt}
%\bibitem{Bland:2013pkt}
  L.~C.~Bland {\it et al.}  [AnDY Collaboration],
  %``Cross Sections and Transverse Single-Spin Asymmetries in Forward Jet Production from Proton Collisions at $\sqrt{s}=500$ GeV,''
  arXiv:1304.1454 [hep-ex].
  %%CITATION = ARXIV:1304.1454;%%
  %3 citations counted in INSPIRE as of 06 Aug 2013

%\cite{Mueller:1999wm}
\bibitem{Mueller:1999wm}
  A.~H.~Mueller,
  %``Parton saturation at small x and in large nuclei,''
  Nucl.\ Phys.\ B {\bf 558}, 285 (1999).
  %%CITATION = HEP-PH/9904404;%%
  %268 citations counted in INSPIRE as of 07 Aug 2013

  %\cite{McLerran:1998nk}
\bibitem{McLerran:1998nk}
  L.~D.~McLerran and R.~Venugopalan,
  %``Fock space distributions, structure functions, higher twists and small x,''
  Phys.\ Rev.\ D {\bf 59}, 094002 (1999).
  %%CITATION = HEP-PH/9809427;%%
  %188 citations counted in INSPIRE as of 07 Aug 2013

  %\cite{Balitsky:2010ze}
\bibitem{Balitsky:2010ze}
  I.~Balitsky and G.~A.~Chirilli,
  %``Photon impact factor in the next-to-leading order,''
  Phys.\ Rev.\ D {\bf 83}, 031502 (2011).
  %%CITATION = ARXIV:1009.4729;%%
  %27 citations counted in INSPIRE as of 07 Aug 2013

  %\cite{Beuf:2011xd}
\bibitem{Beuf:2011xd}
  G.~Beuf,
  %``NLO corrections for the dipole factorization of DIS structure functions at low x,''
  Phys.\ Rev.\ D {\bf 85}, 034039 (2012).
  %%CITATION = ARXIV:1112.4501;%%
  %8 citations counted in INSPIRE as of 07 Aug 2013

%\cite{Kang:2008qh}
\bibitem{Kang:2008qh}
  Z.~-B.~Kang and J.~-W.~Qiu,
  %``Single transverse-spin asymmetry for D-meson production in semi-inclusive deep inelastic scattering,''
  Phys.\ Rev.\ D {\bf 78}, 034005 (2008).
  %%CITATION = ARXIV:0806.1970;%%
  %40 citations counted in INSPIRE as of 07 Aug 2013

  %\cite{Kang:2008ih}
\bibitem{Kang:2008ih}
  Z.~-B.~Kang, J.~-W.~Qiu, W.~Vogelsang and F.~Yuan,
  %``Accessing tri-gluon correlations in the nucleon via the single spin asymmetry in open charm production,''
  Phys.\ Rev.\ D {\bf 78}, 114013 (2008).
    %%CITATION = ARXIV:0810.3333;%%
  %44 citations counted in INSPIRE as of 07 Aug 2013


%\cite{Koike:2011mb}
\bibitem{Koike:2011mb}
  Y.~Koike and S.~Yoshida,
  %``Probing the three-gluon correlation functions by the single spin asymmetry in p^\uparrow p\to DX,''
  Phys.\ Rev.\ D {\bf 84}, 014026 (2011).
  %%CITATION = ARXIV:1104.3943;%%
  %14 citations counted in INSPIRE as of 07 Aug 2013

  %\cite{Koike:2011nx}
\bibitem{Koike:2011nx}
  Y.~Koike and S.~Yoshida,
  %``Three-gluon contribution to the single spin asymmetry in Drell-Yan and direct-photon processes,''
  Phys.\ Rev.\ D {\bf 85}, 034030 (2012).
    %%CITATION = ARXIV:1112.1161;%%
  %7 citations counted in INSPIRE as of 07 Aug 2013





\end{thebibliography}

\documentclass[11pt]{article}
\usepackage{graphicx}
\usepackage{cite}

\begin{document}

%
% Title, Authors, Affiliations
% ===================
%
\title{\Large\bf
Transverse single spin asymmetries at small x
 and the anomalous magnetic moment }

\author{Jian Zhou
 \\[0.3cm]
{\normalsize\it Institut f\"{u}r Theoretische Physik,Universit\"{a}t
Regensburg, Regensburg, Germany}}

\maketitle

\begin{abstract}
\noindent We show that in the Mclerran-Venugopalan model, an axial asymmetrical valence quark distributions in the
transverse plane of a transversely polarized proton can give rise to a spin dependent odderon.
Such polarized odderon is responsible for the transverse single spin asymmetries(SSAs) for jet production
in the backward region of pp collisions and open charm production in semi-inclusive DIS process.
\end{abstract}

\section{Introduction }

The exploration of transverse single spin asymmetries(SSAs)
in high energy scattering experiments has a long history, starting from the mid 1970¡¯s~\cite{Bunce:1976yb}.
The large size of the observed SSAs for single inclusive hadron production
 came as a big surprise and, a priori, posed a challenge
for QCD, as the naive parton model predicts the asymmetries are proportional to
the quark mass~\cite{Kane:1978nd,Ma:2008gm} and thus very small.
During the past few decades,  the remarkable theoretical progress was achieved
by going beyond the naive parton model and following mainly two approaches: one approach is based on transverse momentum dependent(TMD)
factorization~\cite{Sivers:1990fh, Collins:1992kk}
and the other on collinear twist-3
factorization~\cite{Efremov:1981sh,Qiu:1991pp,Ji:1992eu,Yuan:2009dw,Beppu:2010qn}.
In TMD factorization,  naive time reversal odd TMD distributions and fragmentation function,
known as the quark/gluon Sivers functions~\cite{Sivers:1990fh} and the Collins fragmentation function~\cite{Collins:1992kk}
 can account for the large SSAs, while in the collinear twist-3 approach,
the SSAs arise from twist-3 quark gluon correlator
so-called Efremov-Teryaev-Qiu-Sterman function (ETQS)~\cite{Efremov:1981sh,Qiu:1991pp}, tri-gluon
correlation functions~\cite{Ji:1992eu,Beppu:2010qn}, and twist-3 collinear fragmentation functions~\cite{Yuan:2009dw}.

Beyond TMD factorization and collinear twist-3 formalism,
some alternative mechanisms underlying the large SSAs have been proposed, such as
the soft coherent dynamics~\cite{Hoyer:2006hu}, QCD instanton mechanism\cite{Kochelev:1996pv,Qian:2011ya},
and QCD odderon interaction~\cite{Ahmedov:1999ne,Kovchegov:2012ga}.
The authors of papers~\cite{Ahmedov:1999ne} investigated the odderon's contribution
to SSAs in the context of a heavy fermion model. A later calculation formulated in the
saturation/color glass condensate(CGC) framework suggests that SSAs can be generated by the interaction of
spin-dependent light-cone wave function of the projectile with the target gluon field via C-odd odderon exchange~\cite{Kovchegov:2012ga}.
The odderon excitation considered in~\cite{Kovchegov:2012ga} comes from the unpolarized proton/nucleus and is
spin independent. In this paper, however,
 we focus on studying SSAs generated by a spin dependent odderon that comes from the transversely polarized proton.
Such polarized odderon is responsible for SSAs for jet production
in the backward region of polarized proton in pp collisions and open charm production in SIDIS process at small x.

In recent years, the interplay between spin physics and saturation physics is forming into an active field of research.
The early work includes the study of small x evolution of spin dependent structure function $g_1$~\cite{Bartels:1995iu}.
It was pointed out that the spin asymmetries could also be generated by the
 pomeron-odderon interference effect~\cite{oai:arXiv.org:hep-ph/0202231}.
SSAs at forward rapidity in pA collisions have been investigated in Refs.~\cite{Boer:2006rj,Kovchegov:2012ga}.
More recently, the quark/gluon Boer-Mulders distributions inside a large nucleus have been studied
in Refs.~\cite{Metz:2011wb}.

In this paper, we explore  SSA phenomena at small x and identify a spin dependent odderon as the main source of SSAs.
The paper is organized as follows. In Sec.2, we show that, in the MV model, non-vanishing contribution to the spin dependent
 odderon amplitude arises from the left-right asymmetrical color source distribution in the transverse plane of a transversely
polarized proton. In Sec.3, we present compact expressions for SSAs in jet production in the backward region of pp
collisions and open charm production in SIDIS at small x. Both asymmetries are generated by the polarized odderon.
 We summarize our paper in Sec.4.

\section{A spin dependent classical  odderon }
In perturbative QCD, the odderon is a color-singlet exchange
and can be formed by three gluons in a symmetric color state.
It has negative C-parity and therefore dominates the differences between particle-particle and particle-antiparticle scatterings at high energy.
The energy dependence of the odderon exchange is described by the BKP equation~\cite{Bartels:1980pe}.
Within the CGC formalism, one can identify the following operator as the dipole odderon operator~\cite{Hatta:2005as},
\begin{eqnarray}
\hat O(R_\perp,r_\perp)=\frac{1}{2i} \left [ \hat D(R_\perp,r_\perp)-\hat D(R_\perp, -r_\perp) \right ] \ ,
\end{eqnarray}
where
\begin{eqnarray}
\hat D(R_\perp, r_\perp)=\frac{1}{N_c}
{\rm Tr} \left [ U(R_\perp+\frac{r_\perp}{2}) U^\dag(R_\perp-\frac{r_\perp}{2} ) \right ] \ ,
\end{eqnarray}
with the Wilson line being defined as,
\begin{eqnarray}
U(x_\perp)= {\rm P} e^{ig \int_{-\infty}^{+\infty} dx^- A_+(x^-, \ x_\perp)}  \ .
\end{eqnarray}
The small x evolution equation of this odderon operator was constructed using the dipole model\cite{Kovchegov:2003dm}
 and the general JIMWLK equation~\cite{Hatta:2005as}.
In the low parton densities region, the BLV solution~\cite{Bartels:1999yt} to the BKP equation can be recovered from both
the dipole model calculation and the JIMWLK equation with C-odd initial conditions.

The odderon is absent in the original MV model~\cite{McLerran:1993ni} in which the distribution of  large x color source is assumed to take
a Gaussian form. However,  it has been shown that a classical odderon can be generated by
 an additional cubic term  in the modified weight function $W[\rho]$,
 which is given by~\cite{Jeon:2004rk},
\begin{eqnarray}
W[\rho]={\rm exp} \left \{
-\int d^2 x_\perp \left [ \frac{\rho_a(x_\perp) \rho_a(x_\perp)}{2\mu (x_\perp)}
- \frac{g d_{abc} \rho^a(x_\perp) \rho^b(x_\perp)\rho^c(x_\perp) }{4N_c\mu^2(x_\perp)} \right ]
\right \} \ ,
\end{eqnarray}
where $\mu(x_\perp)$is the density of color sources per unit transverse area and related to the valence quark distribution
in the transverse plane: $\int dx_q f_q(x_q, x_\perp)=6  \mu(x_\perp)/g^2$ with $x_q$ being the longitudinal momentum fraction
carried by valence quark;
and $d^{bca}$ is the symmetric structure constant of the color SU(3) group.
Using the above weight function to compute the expectation value of the odderon operator, one obtains~\cite{Jeon:2004rk}
\begin{eqnarray}
&& \int d^2 R_\perp \theta(R_0-|R_\perp|) < \hat O(R_\perp, r_\perp) >
\nonumber\\ &&
=c_0 \alpha_s^3 \int d^2 R_\perp \theta(R_0-|R_\perp|)
\int d^2 z_\perp  {\rm ln}^3 \frac{|R_\perp+r_\perp/2-z_\perp|}{|R_\perp-r_\perp/2-z_\perp|}
e^{-\frac{1}{4}r_\perp^2 Q_s^2} \frac{1}{3} \int d x_q f_q(x_q,z_\perp)
\label{odderon0}
\end{eqnarray}
where $R_0$ is the radius of proton
and the transverse center of  parton longitudinal momentum  is chosen to be as the origin.
The color coefficient $c_0$ is defined as $c_0=\frac{(N_c^2-1)(N_c^2-4)}{4 N_c^3} $.
$Q_s^2 = \alpha_s C_F \mu(R_\perp) {\rm ln} \frac{1}{r_\perp^2 \Lambda_{QCD}^2}$ is the quark saturation momentum.
Here we insert a theta function
 because we have assumed that the dipole must hit the proton directly in order to be able to interact with
quarks inside of it.

To proceed further, we first neglect the dependence of $Q_s^2$ on $R_\perp$ and $r_\perp$. In order to
integrate out $R_\perp$, we use a mathematical trick introduced in~\cite{Kovchegov:2012ga}. One notices that
\begin{eqnarray}
\int d^2 R_\perp{\rm ln}^3 \frac{|R_\perp+r_\perp/2-z_\perp|}{|R_\perp-r_\perp/2-z_\perp|}=0
\end{eqnarray}
if the integration carries over the whole transverse plane.  This result implies,
\begin{eqnarray}
\int d^2 R_\perp{\rm ln}^3 \frac{|R_\perp+r_\perp/2-z_\perp|}{|R_\perp-r_\perp/2-z_\perp|} \theta(R_0-|R_\perp|)
=-\int d^2 R_\perp{\rm ln}^3 \frac{|R_\perp+r_\perp/2-z_\perp|}{|R_\perp-r_\perp/2-z_\perp|}\theta(|R_\perp|-R_0) \ .
\end{eqnarray}
The fact that $|r_\perp|$ is much smaller than $R_0$ for a perturbative dipole  allows us to expand
the integrand on the right-hand-side of the above equation in powers
of $|r_\perp|/|2R_\perp|$ as well as $|z_\perp|/|R_\perp|$. To the first non-trivial order, one has,
\begin{eqnarray}
\int d^2 R_\perp{\rm ln}^3 \frac{|R_\perp+r_\perp/2-z_\perp|}{|R_\perp-r_\perp/2-z_\perp|} \theta(R_0-|R_\perp|)
\approx - \frac{3 \pi}{4 R_0^2} r_\perp^2 (r_\perp \cdot z_\perp)
\label{mathtrick} \ .
\end{eqnarray}
Substituting Eq.~\ref{mathtrick} back into Eq.~\ref{odderon0}, we obtain,
\begin{eqnarray}
\int d^2 R_\perp  \theta(R_0-|R_\perp|) \langle \hat O(R_\perp, r_\perp) \rangle
\approx-\frac{ c_0 \alpha_s^3  \pi}{4 R_0^2} r_\perp^2 e^{-\frac{1}{4}r_\perp^2 Q_s^2}
\int dx_q d^2 z_\perp  (r_\perp \cdot z_\perp)
 f_q(x_q, z_\perp) \ .
 \label{odderon1}
\end{eqnarray}
For a transversely polarized proton, impact parameter dependent valence quark distribution can be parameterized as~\cite{Burkardt:2000za},
\begin{eqnarray}
 f_q(x_q,z_\perp)=  \sum_{u,d} \left \{ {\cal H} (x_q,z_\perp^2)- \frac{1}{2M} \epsilon_{\perp}^{ij}S_{\perp i}
 \frac{\partial{\cal E}(x_q,z_\perp^2)}{ \partial z_{\perp}^j} \right \}
 \label{GPDs} \ ,
\end{eqnarray}
where $S_\perp$ is the proton transverse spin vector and $M$ is the proton mass.
The generalized parton distributions(GPDs) ${\cal H}(x_q,z_\perp^2)$ and ${\cal E}(x_q,z_\perp^2)$ are the
Fourier transformed GPDs $H$ and $E$ with zero skewedness, respectively.
Inserting Eq.~\ref{GPDs} into Eq.~\ref{odderon1}, one immediately obtains
\begin{eqnarray}
\int d^2 R_\perp \theta(R_0-|R_\perp|) < \hat O(R_\perp, r_\perp) >
&=&- \frac{ c_0 \alpha_s^3 \pi  }{8 M R_0^2} e^{-\frac{1}{4}r_\perp^2 Q_s^2}
r_\perp^2 \epsilon_{\perp}^{ij}S_{\perp i}  r_{\perp j}
\int d x_q d^2z_\perp  \sum_{u,d} {\cal E}(x_q,z_\perp^2) \
\nonumber\\
&=&- \frac{ c_0 \alpha_s^3 \pi  }{8 M R_0^2} e^{-\frac{1}{4}r_\perp^2 Q_s^2}
r_\perp^2 \epsilon_{\perp}^{ij}S_{\perp i}  r_{\perp j}
\left ( \kappa_p^u + \kappa_p^d   \right ) \ ,
\label{odderon2}
\end{eqnarray}
where $ \kappa_p^u $ and $ \kappa_p^d $ are the contributions from up and down quarks  to the anomalous magnetic moment
of the proton, respectively. An earlier attempt to connect SSA phenomena to GPD E has been made in paper~\cite{Burkardt:2002ks}.

A few comments are in order on the above analytic result.
\begin{itemize}
\item
First, the odderon exchange under our consideration is clearly  spin dependent.
\item
The polarized odderon originates from the transverse distortion of the
impact parameter dependent PDF inside a transversely polarized proton.
Such transverse distortion has been clearly seen in a lattice QCD calculation~\cite{Gockeler:2006zu}.
\item
Given the nucleon's magnetic moment  $ \kappa_p=1.793$ and $ \kappa_n=-1.913$,
 $ \kappa_p^u $ and $ \kappa_p^d $ can be roughly estimated as $ \kappa_p^u=1.673 $ and $ \kappa_p^d =-2.033 $
 by using the isospin symmetry.
 Obviously, the contributions from u and d quarks to the polarized odderon   largely cancel out.
\end{itemize}

We conclude this section by  making a final remark on our result.
The MV model is expected to work better for a large nucleus.
However, its application to a proton target turns out to be quite successful phenomenologically~\cite{Albacete:2010sy}.
Therefore, qualitatively speaking, our analysis presented here might be relevant in phenomenological studies as well.

\section{Observables }
As discussed in the introduction, the spin dependent odderon is responsible for the transverse single spin asymmetry for
jet production in the backward region of the polarized proton in pp collisions.  The leading order result for jet production
was first derived in Ref.~\cite{Dumitru:2002qt}. Recently, the NLO correction to the cross section has also been calculated
in papers~\cite{Chirilli:2011km}.

At leading order, for the quark-initiated subprocess, the cross section reads
\begin{eqnarray}
\frac{d \sigma^{pA \longrightarrow qX}}{d^2 k_\perp dY}
= \sum_f x q_f(x)
\int \frac{d^2r_\perp }{(2\pi)^2} e^{-i  k_\perp \cdot r_\perp} \int d^2 R_\perp
\langle \hat D(R_\perp,r_\perp) \rangle_{x_g} \ ,
\end{eqnarray}
where $x=\frac{|k_\perp| }{\sqrt{s}}e^{-Y}$ and $x_g=\frac{|k_\perp| }{\sqrt{s}}e^{Y}$, with $Y$ being the rapidity.
$q_f(x)$ is the normal integrated quark distribution from the  unpolarized proton.
Note that we have neglected elastic scattering contribution to the cross section in the above expression.
The dipole S-matrix can be decomposed into the even and odd pieces under the exchange of the transverse coordinates
\begin{eqnarray}
\hat D(R_\perp,r_\perp)=\hat S(R_\perp,r_\perp)+i \hat O(R_\perp,r_\perp) \ ,
\end{eqnarray}
with the symmetric part being defined as
\begin{eqnarray}
\hat S(R_\perp,r_\perp)=\frac{1}{2} \left [ \hat D(R_\perp,r_\perp)+\hat D(R_\perp,-r_\perp) \right ] \ ,
\end{eqnarray}
The cross section then can be re-expressed as
\begin{eqnarray}
\frac{d \sigma^{pA \longrightarrow qX}}{d^2 k_\perp dY}&=&\sum_f x q_f(x)
\int \frac{d^2r_\perp }{(2\pi)^2} e^{-i  k_\perp \cdot r_\perp} \int d^2 R_\perp
\langle \hat S(R_\perp, r_\perp)+i \hat O(R_\perp, r_\perp) \rangle_{x_g}
\nonumber\\ &=&
\sum_f x q_f(x)\left \{  F_{x_g}(k_\perp^2)+\frac{1}{M}
 \epsilon_{\perp}^{ij}S_{\perp i}  k_{\perp j} O_{1T, x_g}^\perp(k_\perp^2) \right \} \ .
\end{eqnarray}
Here we introduce a spin dependent odderon in momentum space: $O_{1T, x_g}^\perp(k_\perp^2)$.
 To some extent, $O_{1T, x_g}^\perp(k_\perp^2)$ can be considered as a C-odd partner of the gluon Sivers function.
In the MV model, the unpolarized gluon distribution is given by,
\begin{eqnarray}
  F_{x_g}(k_\perp^2)=\pi R_0^2 \int \frac{d^2r_\perp }{(2\pi)^2} e^{-i  k_\perp \cdot r_\perp}
e^{-\frac{1}{4} r_\perp^2 Q_{s}^2} \ .
\label{unpolarized}
\end{eqnarray}
Using Eq.~\ref{odderon2}, it is easy to derive,
\begin{eqnarray}
O_{1T, x_g}^\perp(k_\perp^2)=
\frac{ -c_0 \alpha_s^3 \left ( \kappa_p^u + \kappa_p^d   \right )  }{4 R_0^4} \left [
\frac{\partial}{\partial k_\perp^2 } \frac{\partial}{\partial k_\perp^i} \frac{\partial}{\partial k_{\perp i}}
F_{x_g}(k_\perp^2) \right ] \ .
\label{polarized}
\end{eqnarray}
A few comments are in order on the above analytic result.
\begin{itemize}
\item
We note that  $\int d^2 k_\perp k^2_\perp O_{1T, x_g}^\perp(k_\perp^2)=0$.
This relation implies that $O_{1T, x_g}^\perp(k_\perp^2)$ has a node in $k_\perp^2$,
 and the $k_\perp^i$ weighted cross section $\int d^2k_\perp <k_\perp^i d \sigma >$ is zero.
\item
The single spin asymmetry is determined by the ratio $k_\perp O_{1T, x_g}^\perp/F_{x_g}$.
From Eq.~[\ref{unpolarized}] and Eq.~[\ref{polarized}], one finds that
this ratio scales as $k_\perp$ at low transverse momentum, while it scales as $1/k_\perp^3$ at high transverse momentum.
\item
The ratio $k_\perp O_{1T, x_g}^\perp/F_{x_g}$ should drop with decreasing $x_g$ as the power of $(x_g)^{0.3}$ since the leading high energy odderon intercept
is equal to 1 according to the BLV solution.
\end{itemize}
For the gluon initiated channel, the cross section reads
\begin{eqnarray}
\frac{d \sigma^{pA \longrightarrow gX}}{d^2 k_\perp dY}
=  x g(x)
\int \frac{d^2r_\perp }{(2\pi)^2} e^{-i  k_\perp \cdot r_\perp} \int d^2 R_\perp
\langle \hat { \tilde D}(R_\perp,r_\perp) \rangle_{x_g} \ ,
\end{eqnarray}
where $g(x)$ is the normal integrated gluon distribution from the unpolarized proton,
 and $\hat { \tilde D}(r_\perp)$ is given by,
\begin{eqnarray}
\hat {\tilde D}(R_\perp, r_\perp)=\frac{1}{N_c^2-1}
{\rm Tr} \left [ \tilde U(R_\perp+\frac{r_\perp}{2}) \tilde U^\dag(R_\perp-\frac{r_\perp}{2} ) \right ]
\end{eqnarray}
with $\tilde U$ being the Wilson line in the adjoint representation.
In the large $N_c$ limit, the cross section is approximated as,
\begin{eqnarray}
 \frac{d \sigma^{pA \longrightarrow gX}}{d^2 k_\perp dY}
\approx
x g(x)
\int \frac{d^2r_\perp }{(2\pi)^2} e^{-i  k_\perp \cdot r_\perp} \int d^2 R_\perp
{ \Big \langle} \left  \{[ \hat S(R_\perp, r_\perp)]^2+[ \hat O(R_\perp, r_\perp)]^2 \right  \} {\Big \rangle}_{x_g}
 \end{eqnarray}
where all antisymmetric interference terms $\hat S(R_\perp, r_\perp)\hat O(R_\perp, r_\perp)$ completely cancel out.
We are only left with the symmetric terms that do not contribute to the spin asymmetry.
Since SSA vanishes in the gluon initiated jet production process, one might expect that the spin asymmetry
rises with the increasing $|x_F|$ in the backward region of pp collisions.

Recently, the single spin asymmetry in inclusive jet production has been measured in both forward and backward regions
at the AnDY experiment at RHIC~\cite{Nogach:2012sh}.
There is at least one experimental data point in the backward region which is inconsistent
with zero within error bar. The possible two sources of the spin asymmetry in the backward region are the
polarized odderon and the gluon Sivers function. However, we have shown that the gluon Sivers function
dies out very quickly with decreasing $x_g$~\cite{Schafer:2013opa}.
Therefore, this measurement likely indicates that the polarized odderon indeed exists.

Let us now turn to discuss the SSA in open charm production in SIDIS process.
The differential cross section for this process has been calculated in the dipole model~\cite{Mueller:1999wm}
 and in the CGC formalism~\cite{McLerran:1998nk}. The next to leading order correction to this process
 is also available in Refs.~\cite{Balitsky:2010ze,Beuf:2011xd}.
At leading order, the cross section in  momentum space reads,
\begin{eqnarray}
\frac{d \sigma}{dx_B dz  dQ^2 dy  d^2 l_{\perp}}&=&\frac{
\alpha_{em}^2 e_c^2}{2 \pi^4 x_B Q^2} \left [ 1-y+\frac{y^2}{2}\right  ] \left [ z^2+(1-z)^2 \right ]
\nonumber\\&& \times
\int \frac{d^2
p_\perp}{(2\pi)^2} \frac{d^2 k_{\perp}}{(2\pi)^2} \frac{d^2
k_{\perp}'}{(2\pi)^2}
\frac{(l_\perp-k_{\perp})\cdot
(l_\perp-k_\perp')}{[\rho+(l_\perp-k_\perp)^2][\rho+(l_\perp-k_\perp')^2]}
\nonumber\\&&  \times { \Big \langle} {\rm Tr} \Big \{ \left [
U(k_\perp)U^\dag(k_\perp-p_\perp)-(2\pi)^4\delta^2(k_\perp)\delta^2(k_\perp-p_\perp) \right ]
\nonumber\\&&   \ \ \ \ \ \ \ \
  \times \left [
U(k_\perp'-p_\perp) U^\dag(k_\perp')-(2\pi)^4\delta^2(k_\perp')\delta^2(k_\perp'-p_\perp) \right ] \Big \}
{\Big \rangle}_{x_g}
\end{eqnarray}
Here the common kinematical variables in SIDIS process
 are defined as $Q^2=-q \cdot q$, $x_B=Q^2/2P\cdot q$, $y=q\cdot P/P_e \cdot P$
 and $z=l \cdot P/P\cdot q$ where $l$, $P_e$, $P$, and $q$  are momenta for produced charm quark,
 incoming lepton and proton, and virtual photon, respectively.
$\rho$ is defined as $\rho=z(1-z)Q^2$.
$U(k_\perp)$ is the Fourier transform of $U(x_\perp)$.
For simplicity, we have neglected  charm quark mass and only taken into account the
transverse polarized virtual photon contribution to the differential cross section.
 The above formula can be re-organized and expressed in a more compact form:
\begin{eqnarray}
\frac{d \sigma}{dx_B dz  dQ^2 dy  d^2 l_{\perp}}&=&\frac{
\alpha_{em}^2 e_c^2}{2 \pi^4 x_B Q^2}
\int \frac{d^2 k_{\perp}}{(2\pi)^2}
 H(k_\perp,l_\perp,Q^2)
{\Big \langle} {\rm Tr} \left [
U(k_\perp)U^\dag(k_\perp) \right ] {\Big \rangle}_{x_g}
\nonumber\\&=&  \frac{
\alpha_{em}^2 e_c^2N_c}{2 \pi^4 x_B Q^2}
 \int d^2 k_{\perp} H(k_\perp,l_\perp,Q^2)
\left [  F_{x_g}(k_\perp^2)+
 \frac{1}{M} \epsilon_{\perp}^{ ij}S_{\perp i}  k_{\perp j} O_{1T, x_g}^\perp(k_\perp^2) \right ] \ ,
\end{eqnarray}
where the first term recovers the known unpolarized differential cross section,
 whereas the second term is the spin dependent contribution.
For anti-charm quark production, $U(k_\perp)U^\dag(k_\perp)$ that appears in the first line of the above equation should be replaced with
$U^\dag(k_\perp)U(k_\perp)$, leading to the exactly opposite SSA as compared to that in charm quark production.
The hard part $H(k_\perp,l_\perp,Q^2)$ is given by,
\begin{eqnarray}
H(k_\perp,l_\perp,Q^2)=\left [ 1-y+\frac{y^2}{2} \right ] \left [ z^2+(1-z)^2 \right ]
\left [ \frac{l_\perp-k_\perp}{\rho+(l_\perp-k_\perp)^2}-
\frac{l_\perp}{\rho+l_\perp^2} \right ]^2  \ .
\end{eqnarray}

The SSA in open charm production in SIDIS process has also been
calculated using the collinear twist-3 approach~\cite{Beppu:2010qn}(for earlier work, see~\cite{Kang:2008qh}).
In the framework of the collinear factorization, a C-odd tri-gluon
correlation which gives rise to SSA is defined as~\cite{Ji:1992eu,Beppu:2010qn},
\begin{eqnarray}
O^{\alpha \beta \gamma}(x_1,x_2)=
-gi^3\int \frac{dy^- dz^-}{(2\pi)^2 P^+} e^{iy^- x_1P^+}e^{iz^-(x_2- x_1)P^+}
\langle pS| d^{bca} F_b^{\beta+}(0)F_c^{\gamma +}(z^-)F_a^{\alpha+}(y^-)|pS \rangle
\label{tri-gluon}
\end{eqnarray}
where we regard all the free Lorentz indices $\alpha$, $\beta$ and $\gamma$ to be transverse in three dimension.
One can also define a C-even tri-gluon correlation $N^{\alpha \beta \gamma}(x_1,x_2)$ by replacing
$d^{bca}$ with the anti-symmetric tensor $if^{bca}$ in the above equation~\cite{Ji:1992eu,Beppu:2010qn}.
Both the C-even and C-odd tri-gluon correlations contribute to SSAs. However, only the C-odd tri-gluon correlation
is the relevant one at small x as shown in~\cite{Schafer:2013opa}.

It is known that the $k_\perp$ moment of the gluon Sivers function can be related to the gluonic pole C-even
tri-gluon correlation $N^{\alpha \beta \gamma}(x_g,x_g)$. A similar relation between the $k_\perp$ moment of the
polarized odderon and the C-odd tri-gluon correlation can be established.
At small x, exponentials that appear in Eq.~\ref{tri-gluon} can be approximated as
 $e^{iy^- x_1P^+} \approx1 $ and $e^{iz^-(x_2- x_1)P^+}\approx 1$.
In this approximation, one has
\begin{eqnarray}
 \int d^2 k_\perp k_\perp^\alpha k_\perp^\beta k_\perp^\gamma
\frac{1}{M} \epsilon_{\perp}^{ij}S_{\perp i}  k_{\perp j} O_{1T, x_g}^\perp(k_\perp^2)
= \frac{-i g^2 \pi^2}{2 N_c }O^{\alpha \beta \gamma}(x_g) \ ,
\end{eqnarray}
where $O^{\alpha \beta \gamma}(x_g) \equiv O^{\alpha \beta \gamma}(x_1,x_2)$ with $x_g$ being the
total momentum transfer carried by gluons,
which can be conveniently chosen to be $x_g\equiv {\rm Max} \{ x_1,x_2  \}$.

\section{Summary}
In this paper, we have shown that an axial asymmetric color source distribution in the
transverse plane of transversely polarized proton can give rise to a spin dependent odderon in the MV model.
Such polarized odderon is responsible for SSA in jet production in the backward region of pp collisions
and SSA in open charm production in SIDIS process.
As a result, the BLV odderon solution can be tested by studying the x dependence of SSAs.
A relation between $k_\perp$ momentum of the odderon and the collinear twist-3 C-odd tri-gluon correlation
 has been established.
It is straightforward to extend our formalism  to study
SSAs  in open charm production in pp collisions and in Drell-Yan/direct photon processes,
 which already have been calculated in the collinear twist-3 framework~\cite{Koike:2011mb,Koike:2011nx}
 (for earlier work, see~\cite{Kang:2008ih}).
It is  our plan to further explore the possible difference/relation
between the CGC formalism and the collinear twist-3 approach in computing SSAs at small x.

\

\noindent
{\bf Acknowledgments:} I am grateful to Vladimir
Braun for bringing my attention to the potential relation between the odderon and the C-odd tri-gluon correlation.
This work has been supported by BMBF (05P12WRFTE).

\end {document}